OpenStudy (anonymous):

Find f'(x) and f''(x). f(x)=(x^4)(e^x)

OpenStudy (usukidoll):

product rule!

OpenStudy (anonymous):

\[d/dx f(x) = d/dx(x^4)*e^x = d/dx(x^4)*e^x + x^4*d/dx(e^x) = 4x^3*e^x*x^4 = x^4*e*x + 4x^3*e*x\]

OpenStudy (anonymous):

oops that went out of the screen

OpenStudy (usukidoll):

ctrl - to zoom out

OpenStudy (anonymous):

Ah, ok. I attached it also.

OpenStudy (anonymous):

Here is f''(x)

OpenStudy (anonymous):

Thanks! My professor didn't go over this in class and insists on giving problems like this to us for homework.

OpenStudy (anonymous):

I appreciate the help from both of you!

OpenStudy (anonymous):

np